# Safe Payload Transfer with Ship-Mounted Cranes: A Robust Model Predictive Control Approach

### v1

@@ III-A Crane Safety Functions
$\mathcal{B}(t)\!\triangleq\!\big(\bar{\mathbf{p}}_{x}^{l}(t),\bar{\mathbf{p}}_{x}^{u}(t)\big)\!\times\!(\bar{\mathbf{p}}_{y}^{l}(t),\bar{\mathbf{p}}_{y}^{u}(t)\big)\!\times\!(\bar{\mathbf{p}}_{z}^{l}(t),\bar{\mathbf{p}}_{z}^{u}(t)\big).$
@@ III-C Safety-Critical Model Predictive Control
$\begin{array}[]{l}\!{{{\mathbf{x}}^{*}(t)},{\mathbf{u}^{*}(t)}=\ }\displaystyle\!\operatorname*{arg\,min}_{\mathbf{x}(\cdot),\ \mathbf{u}(\cdot)}\!\!~{\mathcal{J}_{T}(t,\!\mathbf{x}(T_{f}))\!+\!\!\!\sum_{t=0}^{T_{f}-1}\!\!\mathcal{J}(t,\!\mathbf{x}(t),\!\mathbf{u}(t))}\\[11.38109pt] ~~~~~~~~~~~~~~~~~\!\textrm{s.t.}\\[5.69054pt] \!\dot{\mathbf{x}}(t)\!=\!\mathbf{f}(t,\mathbf{x}(t))\!+\!\mathbf{g}(\mathbf{x}(t))\mathbf{u}(t),\ \mathbf{x}(0)\!=\!\mathbf{x}_{0},\ \forall t\!\in\![0,T_{f}]\\[5.69054pt] \!h_{\mathbf{t}}(t\!+\!T,\mathbf{F}(t,\mathbf{x}(t),\mathbf{u}(t)))\!-\!h_{\mathbf{t}}(t,\mathbf{x}(t))\!-\!\mathbf{\delta}_{\mathbf{t}}(t,\mathbf{x}(t))\\[2.84526pt] ~~~~~~~~~~~~~~~~~~~~~~~~~\!\geq\!-\alpha(h_{\mathbf{t}}(t,\mathbf{x}(t))),\ \forall t\!\in\![0,T_{f}-1]\\[5.69054pt] \!{h}_{i}(t,\mathbf{x}(t))\!\geq\!0,i\!\in\!\{1,\ldots,6\},\ \forall t\!\in\![0,T_{f}-1]\\[5.69054pt] \!\mathbf{u}(t)\!\in\!\mathcal{U},\ \forall t\!\in\![0,T_{f}-1]\\[5.69054pt] \!\mathbf{x}(t)\!\in\!\mathcal{X},\ \forall t\!\in\![0,T_{f}],\end{array}$
@@
$\begin{array}[]{l}\mathbf{\delta}_{\mathbf{t}}(t,{\mathbf{x}})=\displaystyle\!\min_{\mathbf{\delta}_{\mathbf{t}}\in\mathbb{R}_{\geq 0}}\!\!~~~\mathbf{\delta}_{\mathbf{t}}\\[5.69054pt] ~~~~~~~~~~~~\textrm{s.t.}\\ \mathbf{u}(t)\!\in\!\mathcal{U},\ \!{\mathbf{x}}(t)\!\in\!\mathcal{X}\\[5.69054pt] h_{\mathbf{t}}(t\!+\!T,\mathbf{F}(t,\!{\mathbf{x}}(t),\!\mathbf{u}(t))\!)\!-\!h_{\mathbf{t}}(t_{k}\!+\!\tau,\!\Phi_{\tau}(t,\!{\mathbf{x}},\!\mathbf{u},\mathbf{\delta}_{e})\!)\!-\!\mathbf{\delta}_{\mathbf{t}}\\[2.84526pt] ~~~~~~~~~\geq-\alpha(h_{\mathbf{t}}(t,{\mathbf{x}}(t))),~~\forall\tau\!\in\![0,~T],~~\forall\mathbf{\delta}_{e}\in\mathcal{D}(t).\end{array}$
@@
$\begin{array}[]{l}\mathbf{\delta}_{\mathbf{t}}(t,{\mathbf{x}})=\displaystyle\!\min_{\mathbf{\delta}_{\mathbf{t}}\in\mathbb{R}_{\geq 0}}\!\!~~~\mathbf{\delta}_{\mathbf{t}}\\[2.84526pt] ~~~~~~~~~~~~~~~~~\!\textrm{s.t.}~~~\!\tilde{\mathbf{x}}(t)\!\in\!\{\tilde{\mathbf{x}}_{1},\dots,\tilde{\mathbf{x}}_{N}\},\tilde{\mathbf{x}}_{i}\!\in\!B_{\Delta}({\mathbf{x}})\\[5.69054pt] \!h_{\mathbf{t}}(t\!+\!T,\mathbf{F}(t,{\mathbf{x}}(t),\mathbf{u}^{*}_{k-1}))\!-\!h_{\mathbf{t}}(t\!+\!\tau,\!\mathbf{F}_{\tau}(t,\!\tilde{\mathbf{x}},\!\mathbf{u}^{*}_{k-1}))\!-\!\mathbf{\delta}_{\mathbf{t}}\\[2.84526pt] ~~~~~~~~~~~~~~~~~~~~\geq-\alpha(h_{\mathbf{t}}(t,{\mathbf{x}}(t))),\ \tau\!\in\!\{0,\Delta_{\tau},\ldots,T\}.\end{array}$
We remark that this optimization problem has a single decision variable, and both $\tilde{\mathbf{x}}\}$ and $\tau$ are taken from finite grids; the program simplifies to a one-dimensional linear check: evaluate each sampled constraint and select the most restrictive bound.

### v2

@@ III-A Crane Safety Functions
$\mathcal{B}(t)\!\triangleq\!\big[\bar{\mathbf{p}}_{x}^{l}(t),\bar{\mathbf{p}}_{x}^{u}(t)\big]\!\times\![\bar{\mathbf{p}}_{y}^{l}(t),\bar{\mathbf{p}}_{y}^{u}(t)\big]\!\times\![\bar{\mathbf{p}}_{z}^{l}(t),\bar{\mathbf{p}}_{z}^{u}(t)\big].$
@@ III-C Safety-Critical Model Predictive Control
$\begin{array}[]{l}\!\!{{{\mathbf{x}}^{*}(t)},\!{\mathbf{u}^{*}(t)}\!=\!\ }\displaystyle\!\operatorname*{arg\,min}_{\mathbf{x}(\cdot),\ \mathbf{u}(\cdot)}\!\!~{\mathcal{J}_{T}(T_{f},\!\mathbf{x}(T_{f})\!)\!+\!\!\!\int_{t=0}^{T_{f}}\!\!\!\mathcal{J}(t,\!\mathbf{x}(t),\!\mathbf{u}(t))dt}\\[11.38109pt] ~~~~~~~~~~~~~~~~~\!\textrm{s.t.}\\[5.69054pt] \!\dot{\mathbf{x}}(t)\!=\!\mathbf{f}(t,\mathbf{x}(t))\!+\!\mathbf{g}(\mathbf{x}(t))\mathbf{u}(t),\ \mathbf{x}(0)\!=\!\mathbf{x}_{0},\ \forall t\!\in\![0,T_{f}]\\[5.69054pt] \!h_{\mathbf{t}}(t\!+\!T,\mathbf{F}(t,\mathbf{x}(t),\mathbf{u}(t)))\!-\!h_{\mathbf{t}}(t,\mathbf{x}(t))\!-\!\mathbf{\delta}_{\mathbf{t}}(t,\mathbf{x}(t))\\[2.84526pt] ~~~~~~~~~~~~~~~~~~~~~~~~~\!\geq\!-\alpha(h_{\mathbf{t}}(t,\mathbf{x}(t))),\ \forall t\!\in\![0,T_{f}\!-\!T]\\[5.69054pt] \!{h}_{i}(t,\mathbf{x}(t))\!\geq\!0,i\!\in\!\{1,\ldots,6\},\ \forall t\!\in\![0,T_{f}]\\[5.69054pt] \!\mathbf{u}(t)\!\in\!\mathcal{U},\ \forall t\!\in\![0,T_{f}]\\[5.69054pt] \!\mathbf{x}(t)\!\in\!\mathcal{X},\ \forall t\!\in\![0,T_{f}],\end{array}$
@@
$\begin{array}[]{l}\mathbf{\delta}_{\mathbf{t}}(t,{\mathbf{x}})=\displaystyle\!\operatorname*{arg\,min}_{\mathbf{\delta}_{\mathbf{t}}\in\mathbb{R}_{\geq 0}}\!\!~~~\mathbf{\delta}_{\mathbf{t}}\\[5.69054pt] ~~~~~~~~~~~~\textrm{s.t.}\\ \mathbf{u}(t)\!\in\!\mathcal{U},\ \!{\mathbf{x}}(t)\!\in\!\mathcal{X}\\[5.69054pt] h_{\mathbf{t}}(t\!+\!T,\mathbf{F}(t,\!{\mathbf{x}}(t),\!\mathbf{u}(t))\!)\!-\!h_{\mathbf{t}}(t_{k}\!+\!\tau,\!\Phi_{\tau}(t,\!{\mathbf{x}},\!\mathbf{u},\mathbf{\delta}_{e})\!)\!-\!\mathbf{\delta}_{\mathbf{t}}\\[2.84526pt] ~~~~~~~~~\geq-\alpha(h_{\mathbf{t}}(t,{\mathbf{x}}(t))),~~\forall\tau\!\in\![0,~T],~~\forall\mathbf{\delta}_{e}\in\mathcal{D}(t).\end{array}$
@@
$\begin{array}[]{l}\mathbf{\delta}_{\mathbf{t}}(t,{\mathbf{x}})=\displaystyle\!\operatorname*{arg\,min}_{\mathbf{\delta}_{\mathbf{t}}\in\mathbb{R}_{\geq 0}}\!\!~~~\mathbf{\delta}_{\mathbf{t}}\\[2.84526pt] ~~~~~~~~~~~~~~~~~\!\textrm{s.t.}~~~\!\tilde{\mathbf{x}}(t)\!\in\!\{\tilde{\mathbf{x}}_{1},\dots,\tilde{\mathbf{x}}_{N}\},\tilde{\mathbf{x}}_{i}\!\in\!B_{\Delta}({\mathbf{x}})\\[5.69054pt] \!h_{\mathbf{t}}(t\!+\!T,\mathbf{F}(t,{\mathbf{x}}(t),\mathbf{u}^{*}_{k-1}))\!-\!h_{\mathbf{t}}(t\!+\!\tau,\!\mathbf{F}_{\tau}(t,\!\tilde{\mathbf{x}},\!\mathbf{u}^{*}_{k-1}))\!-\!\mathbf{\delta}_{\mathbf{t}}\\[2.84526pt] ~~~~~~~~~~~~~~~~~~~~\geq-\alpha(h_{\mathbf{t}}(t,{\mathbf{x}}(t))),\ \tau\!\in\!\{0,\Delta_{\tau},\ldots,T\}.\end{array}$
We remark that this optimization problem has a single decision variable, and both $\tilde{\mathbf{x}}$ and $\tau$ are taken from finite grids; the program simplifies to a one-dimensional linear check: evaluate each sampled constraint and select the most restrictive bound.